DIMINISHED CAGE EFFECT IN p-H_{2}: INFRARED SPECTRA OF CH_{3}S OBSERVED FROM PHOTOLYSIS OF CH_{3}SH, CH_{3}SCH_{3}, AND CH_{3}SSCH_{3} ISOLATED IN p-H_{2}

\underline{YUAN-PERN LEE}, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan; MOHAMMED BAHOU, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan.

We report infrared absorption spectrum of the methylthio (or thiomethoxy) radical, CH_{3}S, isolated in solid p-H_{2}. CH_{3}S was produced by in situ UV photodissociation of three precursors: CH_{3}SH, CH_{3}SCH_{3}, and CH_{3}SSCH_{3} isolated in solid p-H_{2}. New absorption features commonly observed with similar intensity ratios in experiments using these precursors are assigned as absorption of CH_{3}S. In addition to the previously assigned transitions of ν_{3} (a_{1}) at 727.1 cm^{−1}, fundamental transitions ν_{6} (a_{1}) at 771.1, ν_{6} (e) at 1056.6, ν_{5} (a_{1}) at 1400.0, and ν_{4} (a_{1}) at 2898.0 cm^{−1} were observed. The wavenumbers of these features agree satisfactorily with those predicted with a spin-vibronic Hamiltonian accounting for the anharmonic effects and the Jahn-Teller effects to the quartic term;^{a} the corresponding wavenumbers predicted from theory are ν_{6} (a_{1}) at 793, ν_{6} (e) at 1105, ν_{5} (a_{1}) at 1436, and ν_{4} (a_{1}) at 2938 cm^{−1}, with deviations of 1.4-4.6 % from experiments. Previous attempts of UV photolysis of CH_{3}SCH_{3} and CH_{3}SSCH_{3} isolated in an Ar matrix failed to produce CH_{3}S. These results serve as an excellent example that the diminished cage effect of solid p-H_{2} makes production of free radicals via photolysis in situ feasible. If time permits, other examples will be discussed.

aA. V. Marenich and J. E. Boggs, J. Chem. Theory Comput. \underline{1}, 1162 (2005).